8. In a particular diffraction pattern, the green component 555 nm) in the fourth-order spectrum is deviated at an angle of 50°. How many lines per millimeter does the grating have? Show my other submissions Ox 118783784 lines/mm
Answer #1

\lambda=555nm=555x10-9 m

m=4

theta=50 degrees

we know that distance between the slits= d= m\lambda/sin theta

d= 4 (555x10-9)/sin50

d= 2.89x10-6 m

d=2.89x10-3 mm

number of lines per mm is N= 1/d

N= 1/(2.89x10-3)

N=345 lines/mm